[Brahms-l] Shift report 20050526 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu May 26 2005 - 15:59:15 EDT
Shifter leader was JMH
Shifters were 


1. Beam for 1/2 hour, had 2 hour access.
2. Made one run
3. no detector problems

Durning access CC HV moved from BB3 to BB2 (RAS)
MRST0 counters taken down (EB)

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 26 15:59:27 2005

This archive was generated by hypermail 2.1.8 : Thu May 26 2005 - 15:59:44 EDT